how do i answer this? using substitution elimination
y+3x=6
y=2x+1

Answer by Alan3354(69443)

how do i answer this? using substitution elimination
y+3x=6
y=2x+1
--------
Using substitution:
y is solved for in the 2nd eqn, so sub for y in the 1st one.
----
(2x+1) + 3x = 6
5x + 1 = 6
5x = 5
x = 1
---------
y = 2x + 1
y = 3